### v7

@@ 4.6 Comparison Experiment
To ensure fairness in the experiments, we used the same training set and data augmentation techniques described in Section 4.3.1 for all the compared models. Unlike Suh et al. [31], who split the input images into $256\times 256$ patches, our proposed method uses $224\times 224$ patches. To demonstrate the positive impact of this change on the model performance, we conducted experiments comparing models using $224\times 224$ patches with those using $256\times 256$ patches, keeping all other settings unchanged.
Table 2 presents the performance of the models across different patch sizes on six DIBCO datasets. As established in Section 4.5, the images processed with DWT and normalization outperform those processed with either the original input images or only with DWT. Consequently, all models in this section utilize images processed with DWT and normalization. According to Table 2, the mean values of FM, p-FM, PSNR, DRD, and Avg for the model based on $224\times 224$ patches are 93.57, 96.83, 20.59, 2.09, and 77.22, respectively, which are superior to those for the model based on $256\times 256$ patches. This indicates that the model based on $224\times 224$ patches performs better, and therefore we will use this model for subsequent comparisons.
@@ 4.7 Experimental Results
Different from the conference paper [51] of this work, which uses option 3 to train the model, the proposed method in this paper applies DWT and normalization to the images in Stage-1 as input to train the model. The performance of the trained model will then be compared with that of other methods.
@@
Figure 7 presents the visualization results at each stage of the trained model. Specifically, five images were randomly selected from the PHIBD and Bickley Diary datasets to demonstrate the step-by-step process of image enhancement and binarization using the proposed method. Figure 7 (b) represents the result of retaining the LL subband images after applying DWT and normalization. This stage performs noise reduction on the original input image, which is used as the ground truth image for the generator. Figure 7 (c) shows the result of image enhancement achieved using adversarial networks, where the image has the background color removed and the color of the text highlighted. Figure 7 (d) is the final output image obtained using the proposed whole method. Notably, it is evident that our final output result closely approximates Figure 7 (e), the ground truth image.
@@
Table 3 presents the evaluation results of different models on six DIBCO datasets. Overall, the proposed method achieves the top two best performance in terms of FM, p-FM, PSNR, and DRD, with its Avg value reaching the SOTA level across all six datasets. Notably, the performance of the proposed method on the DIBCO 2011 and 2013 datasets is particularly good, obtaining the best values for each of the evaluated metrics. For instance, on the DIBCO 2011 dataset, the p-FM value of the proposed method is 97.47, significantly surpassing the second best value of 95.93. Similarly, on the DIBCO 2018 dataset, the FM value of the proposed method is 92.89, considerably higher than the second best value of 91.86. These results indicate that the proposed method is more effective than other methods for the binarization of degraded color documents.
@@
In addition to assessing method performance by comparing quantitative metrics across various methods, this work aims to provide readers with an intuitive understanding of the distinctions by the visualization results. Consequently, we have chosen images from the DIBCO 2011 and 2013 datasets to illustrate examples of binarization results. Figures 8 and 9 show the binarization results of images HW1 and HW5 by different methods. These figures visually demonstrate that the proposed method outperforms deep learning based methods and the traditional binarization methods in terms of the shadow and noise elimination. Furthermore, the proposed method is good at preserving textual content while effectively mitigating the presence of the shadows and noise.
5 Conclusion
@@
This paper introduces a novel three-stage method based on DWT and GANs to apply the task of text degradation in degraded document images with color backgrounds. The proposed method involves preserving the LL subband images through DWT to eliminate image noise, serving as the ground truth images for the generator. It trains independent networks utilizing RGB three single-channel images and one grayscale image, followed by a fusion of local and global prediction networks for binarization. As far as our knowledge extends, this work first represents the three-stage method that combines traditional image processing with deep learning for document binarization. The experimental results demonstrate that employing U-Net++ with EfficientNet-B6 as the generator and PatchGAN as the discriminator outperforms performance compared to other SOTA methods.
@@ 6.3 Declaration of Generative AI and AI-assisted technologies in the writing process
The authors only use generative artificial intelligence (AI) and AI-assisted technologies to improve language.

### v8

@@ 4.6 Comparison Experiment
To ensure fairness in the experiments, we use the same training set and data augmentation techniques described in Section 4.3.1 for all the compared models. Unlike Suh et al. [31], who split the input images into $256\times 256$ patches, our proposed method sets the patch size to $224\times 224$. To demonstrate the positive impact of this change on the model performance, we conduct the experiments comparing models using $224\times 224$ patches with those using $256\times 256$ patches, keeping all other settings unchanged. Table 2 presents the performance of the models across different patch sizes on six DIBCO datasets. Since we have demonstrated in Section 4.5 that the DWT with normalization technique outperforms the only DWT technique, the models in Table 2 all use the DWT with normalization technique.
According to Table 2, the mean values of FM, p-FM, PSNR, DRD, and Avg for the model based on $224\times 224$ patches are 93.57, 96.83, 20.59, 2.09, and 77.22, respectively, which are superior to those for the model based on $256\times 256$ patches. This indicates that the model based on $224\times 224$ patches has better model performance, and therefore we will use the model based on $224\times 224$ patches for subsequent comparisons.
@@ 4.7 Experimental Results
Different from the conference paper [53] of this work that uses the DWT with normalization to only process the input images, this paper uses the DWT with normalization to process both the input images and the Ground-Truth images of the three single-channel images (red, green, and blue) in Stage-1.
@@
Figure 8 presents the visualization results at each stage of the proposed model. Specifically, four images are randomly selected from the PHIBD [47] and BD [48] datasets to demonstrate the step-by-step process of document image enhancement and binarization. Figure 8 (b) represents the results of the LL subband images procossed by the DWT, and (c) represents the results of the LL subband images procossed by the DWT with normalization. It can be seen that there is a significant difference between (b) and (c), which performs noise reduction on the original input images. Figure 8 (d) shows the effect of image enhancement achieved using the GANs models, where the images remove the background color and highlight the color of the text. Figure 8 (e) presents the final output images obtained using the whole proposed method. It is worth noting that the final outputs are very close to Figure 8 (f) the Ground-Truth images.
@@
Table 9(f) presents the evaluation results of different models on six DIBCO datasets. Overall, the proposed method achieves the top two best performance in terms of FM, p-FM, PSNR, and DRD, while its Avg value reaching the SOTA level across all six datasets. Notably, the performances of the proposed method on the DIBCO 2011 [11] and H-DIBCO 2018 [16] datasets are particularly good, obtaining the best values for each of the evaluation metrics. For instance, on the DIBCO 2011 dataset, the p-FM value of the proposed method is 97.47, significantly surpassing the second best value of 95.93. Similarly, on the H-DIBCO 2018 dataset, the FM value of the proposed method is 92.89, considerably higher than the second best value of 91.86. These results indicate that the proposed method is more effective than other methods using the same training set for document image enhancement and binarization.
@@
In addition to assessing method performance by comparing the quantitative metrics of different methods, this work aims to allow the reader to intuitively understand the differences between the methods by visualizing the binarization results. Figures 9 and 10 show the examples the binarization results by different methods for the image “HW1” of the DIBCO 2011 dataset, and “HW5” of the DIBCO 2013 dataset. These figures visually demonstrate that the proposed method outperforms the traditional threshold methods and deep learning-based methods in terms of the shadow and noise elimination. In addition, the proposed method retains the text content well while effectively mitigating the effects of shadows and noise. Specifically, as shown in Figure 10 (g) and (h), the proposed method is more efficient in eliminating noise compared to Suh et al.’s [31] method.
5 Conclusion and Future Work
This paper introduces an effective three-stage network method for enhancement and binarization of gray-scale and color degraded document images. Compared to other methods, the proposed method is more efficient in removing noise from degraded document images, and this effect applies to both gray-scale and color images.
@@ 5 Conclusion and Future Work
As far as our knowledge extends, this work first represents the three-stage network method that combines traditional image processing with deep learning for document image enhancement and binarization. The experimental results show that the performance of the proposed method using U-Net++ with EfficientNet-B6 as generator and PatchGAN as discriminator is better compared to other methods. In addition, based on the experimental results, this paper also demonstrates that for document images, the binarization results of the images processed by the DWT with normalization are closer to the Ground-Truth images than the binarization results of the original input images.
We believe that this paper is of great significance to the reconstruction and analysis of ancient documents. In future work, we will apply the proposed method to more degraded document images from different cultures to extract useful information to help culture organizations to analyze the history of ancient civilizations. In particular, we will focus on the complex structure and polysemous characters, such as Chinese, Egyptian, Arabic, etc.
@@ 6.3 Declaration of Generative AI and AI-assisted technologies in the writing process
The authors only use generative artificial intelligence (AI) and AI-assisted technologies to improve the language.